# Ultrafast All-Optical Measurement of Squeezed Vacuum in a Lithium Niobate Nanophotonic Circuit

### v3

@@ IV.1 Shot-noise Calibration
Data from the shot-noise calibration procedure is show in figure 5A. During calibration, the mechanical shutter is initially left open. As the relative delay of the squeezer path is scanned, the measurement amplifier oscillates between measuring squeezed (with destructive pump interference) and anti-squeezed (with constructive pump interference) quadratures, resulting in the black curve in Fig. 5A. The remaining pump light collected from the output is monitored on a second OSA to determine the strength of the pump interference. The shutter is then closed and the baseline shot noise is measured by the OSA resulting in the gray curve in Fig. 5A. To find the shot-noise level corresponding to destructive pump interference (i.e. the shot-noise minimum), the intensity of the measurement OPA pump is lowered until the measured power at the pump OSA matches the minima of the previously measured pump interference. The shot-noise minimum is then measured on the slow photodetector and OSA resulting in the red curve in Fig. 5A. This same procedure is repeated for the shot-noise maximum to calibrate the anti-squeezing and obtain the blue curve. The pump OSA is set to measure around 970 nm to minimize any pump depletion effects. After shot-noise calibration, the squeezing is calculated as:
@@ IV.2 Slow Photodetector Measurements
Figure 5B shows the measured squeezing on the slow photodetector and OSA as a function of the input pulse energy for the squeezer waveguide with a measurement pulse energy of 30 pJ. For these measurements, the 1950-nm long-pass filter is removed to accentuate the impact of multiple modes (a Schmidt number of 2.8). Because the OSA uses a 2-nm bandpass filter, it can better suppress higher-order modes to isolate the squeezing in the fundamental mode [18, 33]. For the photodetector, the presence of higher-order modes with less squeezing reduces overall measured squeezing as the photodetector is not mode-sensitive and measures the squeezing as averaged over all modes.
@@
Equation 10 also allows us to calculate the expected anti-squeezing, however, the measured data has a stark divergence from the expected trend. This is a result of gain-compression in the measurement OPA. Because the measurement OPA operates in the pump-depleted regime, the anti-squeezed vacuum cannot be sufficiently amplified to reflect the true anti-squeezing entering the OPA. The same is true for the squeezing measurements as the shot-noise level becomes compressed, leading to lower measured squeezing.
@@ IV.3 Pulse-to-pulse Measurements
Figure 6: A) Sample photon number distributions from fast measurements at different $\phi$. Red points are data, blue curve is a 2-mode fit. B) Photon number variance vs $\phi$ for slow and fast (pulse-to-pulse) measurements, and the 1st mode of a 2-mode fit. C) Recovered squeezed vacuum Wigner function. D) Expected density matrix ($\rho$) based on measured squeezing and anti-squeezing. E) Recovered $\rho$. F) Absolute difference between expected and recovered $\rho$.
@@
Fast photodetector measurements resolve the pulse-to-pulse intensity and provide the statistical information necessary to resolve $P(N,\phi)$ and recover the Wigner function. During a fast measurement, the shutter is initially left open, and the fast oscilloscope is triggered using the voltage applied to the piezo driving the squeezer’s delay stage to coincide with the middle of the $\phi$ scan. Data from the slow oscilloscope is used to determine $\phi$ and estimate the coupling stability based on measurement-to-measurement variations in the shot noise. After recording data, the shutter is closed to block the squeezer and the measurement is repeated to find the calibrated shot-noise level.
@@
Figure 6 shows results gathered from fast photodetector measurements for a squeezer pulse energy of 5 pJ and a measurement pulse energy of 45 pJ. This measurement pulse energy was chosen to satisfy SNR requirements at the fast photodetector given the loss between the chip and detector. Sample measured histograms at 4 different phases are plotted in Fig. 6A. Each histogram corresponds to a single data point in Fig. 6B. For each measurement, we fit a 2-mode photon number distribution model to extract the variance of the fundamental mode as a function of $\phi$. Figure 6B shows these measurements for three cases: slow photodetector measurements, fast photodetector measurements, and the 2-mode fit. Slow measurement values are calculated from the slow oscilloscope trace using the method outlined in section IV.2. Because of the photon number offset resulting from pump depletion, the numerator and denominator of Eq.9 have a small but constant offset, resulting in the measured squeezing and anti-squeezing being underestimated. Fast measurement values are calculated using the variance of the photon numbers measured at the fast detector. This overcomes the offset problem and measures more squeezing and anti-squeezing, but these measurements are still contaminated with remaining contributions from higher-order modes. The 2-mode fit addresses this problem by isolating the variance as a function of $\phi$ for the fundamental mode, leading to higher measured squeezing and anti-squeezing. After performing this fit for all measurements, the distributions are re-sampled and fed into a maximum likelihood algorithm which calculates the Wigner function (Fig. 6C). We extract a squeezing in the fundamental mode of 2.41 $\pm$ 0.34 dB and an anti-squeezing of 3.87 $\pm$ 0.61 dB. Figure 6E shows the recovered density matrix corresponding to the Wigner function plotted in Fig. 6C. We show the expected density matrix given the measured squeezing and anti-squeezing in Fig. 6D, and the absolute difference between the expected and recovered in Fig. 6F where we calculate a fidelity of 0.9998 $\pm$ 0.0001. This fidelity serves as a sanity-check to ensure our data analysis is not introducing distortions into the measurement. While our fitting strategy limits us to states with known distributions, we can overcome this limitation with tighter bandpass filtering to better reject higher-order modes. Our losses between the chip and fiber (10 dB) and from our filter setup (6 dB) currently prevent us from tighter filtering as the remaining light is no longer intense enough to provide a sufficient signal-to-noise ratio at the fast photodetector.
@@ V Conclusion
We have demonstrated OPA-based Wigner tomography for squeezed vacuum in the ultrafast regime using dispersion-engineered TFLN. This is, to the extend of the knowledge of the authors, the first demonstration of all-optical quantum state tomography on-chip. The low dispersion of our OPA design allows operation with ultra-short pulses to exceed the bandwidth of traditional measurement techniques and access a new regime of THz repetition-rate measurement and computation. Achieving THz clock rates is a fundamental advancement for time-multiplexed systems such as cluster states and measurement-based quantum computation as higher clock rates translates to faster computation and allows for larger quantum circuits to be realized.
Our THz claim is based on dispersion calculations derived from the waveguide geometry measured using atomic force microscopy. We also note that similar dispersion-engineered OPAs with faster clock rates than the one presented here have been experimentally characterized [40]. Speed limitations imposed by the photodetector bandwidth are present in our experiment, but can be overcome with fast demultimplexing schemes which we elaborate in the supplementary.

### v4

@@ IV.1 Shot-noise Calibration
Data from the shot-noise calibration procedure are show in figure 5A. During calibration, the mechanical shutter is initially left open. As the relative delay of the squeezer path is scanned, the measurement amplifier oscillates between measuring squeezed (with destructive pump interference) and anti-squeezed (with constructive pump interference) quadratures, resulting in the black curve in Fig. 5A. The remaining pump light collected from the output is monitored on a second OSA to determine the strength of the pump interference. The shutter is then closed and the baseline shot noise is measured by the OSA resulting in the gray curve in Fig. 5A. To find the shot-noise level corresponding to destructive pump interference (i.e. the shot-noise minimum), the intensity of the measurement OPA pump is lowered until the measured power at the pump OSA matches the minima of the previously measured pump interference. The shot-noise minimum is then measured on the slow photodetector and OSA resulting in the red curve in Fig. 5A. This same procedure is repeated for the shot-noise maximum to calibrate the anti-squeezing and obtain the blue curve. The pump OSA is set to measure around 970 nm to minimize any pump depletion effects. After shot-noise calibration, the squeezing is calculated as:
@@ IV.2 Slow Photodetector Measurements
Figure 5B shows the measured squeezing on the slow photodetector and OSA as a function of the input pulse energy for the squeezer waveguide with a measurement pulse energy of 30 pJ. For these measurements, the 1950-nm long-pass filter is removed to accentuate the impact of multiple modes (a Schmidt number of 2.8). Because the OSA uses a 2-nm bandpass filter, it can better suppress higher-order modes to isolate the squeezing in the fundamental mode [17, 33]. For the photodetector, the presence of higher-order modes with less squeezing reduces overall measured squeezing as the photodetector is not mode-selective and measures the squeezing as averaged over all modes.
@@
Equation 10 also allows us to calculate the expected anti-squeezing, however, the measured data have a stark divergence from the expected trend. This is a result of gain-compression in the measurement OPA. Because the measurement OPA operates in the pump-depleted regime, the anti-squeezed vacuum cannot be sufficiently amplified to reflect the true anti-squeezing entering the OPA. The same is true for the squeezing measurements as the shot-noise level becomes compressed, leading to lower measured squeezing.
@@ IV.3 Pulse-to-pulse Measurements
Figure 6: A) Sample photon number distributions from fast measurements at different $\phi$. Red points are data, and the blue curve is a two-mode fit. B) Photon number variance vs $\phi$ for slow, fast (pulse-to-pulse) measurements and the first mode of a two-mode fit. Solid lines are the expected cosine dependence taken from Ref.[30]. C) Recovered squeezed vacuum Wigner function. Solid and dashed while lines lie along the $1/(e\pi)$ contours. D) Expected density matrix ($\rho$) based on measured squeezing and anti-squeezing. E) Recovered $\rho$. F) Absolute difference between expected and recovered $\rho$.
@@
Fast photodetector measurements resolve the pulse-to-pulse intensity and provide the statistical information necessary to resolve $P(N,\phi)$ and recover the Wigner function. During a fast measurement, the shutter is initially left open, and the fast oscilloscope is triggered using the voltage applied to the piezo driving the squeezer’s delay stage to coincide with the middle of the $\phi$ scan. Data from the slow oscilloscope are used to determine $\phi$ and estimate the coupling stability based on measurement-to-measurement variations in the shot noise. After recording data, the shutter is closed to block the squeezer and the measurement is repeated to find the calibrated shot-noise level.
@@
Figure 6 shows results gathered from fast photodetector measurements for a squeezer pulse energy of 5 pJ and a measurement pulse energy of 45 pJ. This measurement pulse energy was chosen to satisfy SNR requirements at the fast photodetector given the loss between the chip and detector. Sample measured histograms at four different phases are plotted in Fig. 6A. Each histogram corresponds to a single data point in Fig. 6B. For each measurement, we fit a two-mode photon number distribution model to extract the variance of the fundamental mode as a function of $\phi$. Figure 6B shows these measurements for three cases: slow photodetector measurements, fast photodetector measurements, and the two-mode fit. Slow measurement values are calculated from the slow oscilloscope trace using the method outlined in section IV.2. Because of the photon number offset resulting from pump depletion, the numerator and denominator of Eq.9 have a small but constant offset, resulting in the measured squeezing and anti-squeezing being underestimated. Fast measurement values are calculated using the variance of the photon numbers measured at the fast detector. This overcomes the offset problem and measures more squeezing and anti-squeezing, but these measurements are still contaminated with remaining contributions from higher-order modes. The two-mode fit addresses this problem by isolating the variance as a function of $\phi$ for the fundamental mode, leading to higher measured squeezing and anti-squeezing. After performing this fit for all measurements, the distributions are re-sampled and fed into a maximum likelihood algorithm that calculates the Wigner function (Fig. 6C). We extract a squeezing in the fundamental mode of 2.41 $\pm$ 0.34 dB and an anti-squeezing of 3.87 $\pm$ 0.61 dB. Figure 6E shows the recovered density matrix corresponding to the Wigner function plotted in Fig. 6C. We show the expected density matrix given the measured squeezing and anti-squeezing in Fig. 6D, and the absolute difference between the expected and recovered in Fig. 6F where we calculate a fidelity of 0.9998 $\pm$ 0.0001. This fidelity serves as a sanity-check to ensure our data analysis is not introducing distortions into the measurement. While our fitting strategy limits us to states with known distributions, we can overcome this limitation with tighter bandpass filtering to better reject higher-order modes. Our losses between the chip and fiber (10 dB) and from our filter setup (6 dB) currently prevent us from tighter filtering as the remaining light is no longer intense enough to provide a sufficient signal-to-noise ratio at the fast photodetector.
Figure 7: A diagram of the squeezer and measurement OPA circuit.
@@ V Conclusion
We have demonstrated OPA-based Wigner tomography for squeezed vacuum in the ultrafast regime using dispersion-engineered TFLN. This is, to the best of the authors’ knowledge, the first demonstration of all-optical quantum state tomography on-chip. The low dispersion of our OPA design allows operation with ultra-short pulses to exceed the bandwidth of traditional measurement techniques and access a new regime of THz repetition-rate measurement and computation. Achieving THz clock rates is a fundamental advancement for time-multiplexed systems such as cluster states and measurement-based quantum computation as higher clock rates translate to faster computation and allow for larger quantum circuits to be realized.
Our THz claim is based on dispersion calculations derived from the waveguide geometry measured using atomic force microscopy. We also note that similar dispersion-engineered OPAs with faster clock rates than the one presented here have been experimentally characterized [40]. Speed limitations imposed by the photodetector bandwidth are present in our experiment, but can be overcome with fast demultiplexing schemes which we elaborate in Appendices.
@@ VI Acknowledgments
VII Data Availability Statement
The data that support the findings of this article are not publicly available upon publication because it is not technically feasible and/or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.
Figure 8: A) A plot of the joint-spectral intensity function of the squeezed vacuum produced by our OPA. B) The first four modes of the JSI calculated via the Bloch-Messiah decomposition. C) Plot of the mode number vs proportion of mode present in the JSI.
VIII Appendix 1: Multimode Cascaded Squeezing in the Heisenberg picture
We use the formalism of Ref.[38] to derive the output of two cascaded multimode OPAs: a squeezer OPA and a measurement OPA. Figure 7 shows the circuit diagram. We begin with the state after the squeezer OPA (denoted with superscript $s$) and decompose the annihilation operator into the squeezer OPA’s eigenmodes (denoted by subscript $m$):
$\hat{a}^{s}(t)$ $=\sum_{m}\psi^{s*}_{m}(t)\hat{a}^{s}_{m}$ (11)
$=\sum_{m}\psi^{s*}_{m}(t)(\cosh r^{s}_{m}\hat{a}^{in}_{m}+\sinh r^{s}_{m}\hat{a}^{in{\dagger}}_{m})$
where $r_{m}^{s}$ is the squeezing parameter of mode $m$. Each eigenmode of the squeezer OPA is independently squeezed. To simplify later expressions, we assume that $r_{m}^{s}=r_{m}^{s}(\phi)$ where $\phi$ is the phase of the squeezed vacuum relative to the measurement pump. In our experiment, we inject vacuum into the squeezer OPA, and so $\hat{a}^{in}\equiv\hat{a}^{vac}$.
After the squeezer OPA, the state is placed into the measurement OPA (superscript $ms$ and mode subscript $n$) and the resulting output operator is labeled $\hat{a}^{out}(t)$. The measurement OPA amplifies its own modes as independent single-mode OPAs:
$\hat{a}^{out}(t)$ $=\sum_{n}\psi^{ms*}_{n}(t)\hat{a}^{out}_{n}$ (12)
$=\sum_{n}\psi^{ms*}_{n}(t)(\cosh r^{ms}_{n}\hat{a}^{s}_{n}+\sinh r^{ms}_{n}\hat{a}^{s{\dagger}}_{n})$
We can write the squeezed operators in the measurement OPA basis as a function of the mode overlaps between the measurement OPA’s and squeezer OPA’s eigenmodes $\sigma_{mn}=\int dt~\psi^{ms}_{n}(t)\psi^{s*}_{m}(t)$ :
$\hat{a}^{s}_{n}$ $=\int_{-\infty}^{\infty}dt~\psi^{ms}_{n}(t)\hat{a}^{s}(t)$ (13)
$=\int_{-\infty}^{\infty}dt~\psi^{ms}_{n}(t)\sum_{m}\psi^{s*}_{m}(t)\hat{a}^{s}_{m}$
$=\sum_{m}\sigma_{mn}\hat{a}^{s}_{m}.$
The output annihilation operator is then:
$\hat{a}^{out}(t)$ $=\sum_{mn}\psi^{ms*}_{n}(t)(\cosh r^{ms}_{n}\sigma_{mn}\hat{a}^{s}_{m}+\sinh r^{ms}_{n}\sigma^{*}_{mn}\hat{a}^{s{\dagger}}_{m})$ (14)
$=\sum_{mn}\psi^{ms*}_{n}(t)$
$(\cosh r^{ms}_{n}\sigma_{mn}(\cosh r^{s}_{m}\hat{a}^{in}_{m}+\sinh r^{s}_{m}\hat{a}^{in{\dagger}}_{m})$
$~~~~~~~~+\sinh r^{ms}_{n}\sigma^{*}_{mn}(\cosh r^{s}_{m}\hat{a}^{in{\dagger}}_{m}+\sinh r^{s}_{m}\hat{a}^{in}_{m}))$
For large measurement OPA gains, we use the approximation $\cosh r^{ms}_{n}\approx\sinh r^{ms}_{n}\approx\frac{1}{2}\exp(r^{ms}_{n})$, when $e^{r^{ms}}\gg 1\gg e^{-r^{ms}}$. Then, the output operator becomes:
$\hat{a}^{out}(t)=\sum_{mn}\frac{e^{r^{ms}_{n}}}{2}\psi^{ms*}_{n}(t)$
$(\sigma_{mn}(\cosh r^{s}_{m}\hat{a}^{in}_{m}+\sinh r^{s}_{m}\hat{a}^{in{\dagger}}_{m})+$
$\sigma^{*}_{mn}(\cosh r^{s}_{m}\hat{a}^{in{\dagger}}_{m}+\sinh r^{s}_{m}\hat{a}^{in}_{m}))$ (15)
We can write the photon number operator $\hat{N}$ for an individual pulse as:
$\hat{N}$ $=\int_{-\Delta t}^{\Delta t}dt~\hat{a}^{out{\dagger}}(t)\hat{a}^{out}(t)$ (16)
$=\int dt\sum_{m,m^{\prime},n,n^{\prime}}\frac{e^{r^{ms}_{n}+r^{ms}_{n^{\prime}}}}{4}$
$\Big(\psi^{ms}_{n}(t)(\sigma^{*}_{mn}(\cosh r^{s}_{m}\hat{a}^{in{\dagger}}_{m}+\sinh r^{s}_{m}\hat{a}^{in}_{m})$
$~~~~~~~+\sigma_{mn}(\cosh r^{s}_{m}\hat{a}^{in}_{m}+\sinh r^{s}_{m}\hat{a}^{in{\dagger}}_{m}))\Big)$
$~~~~\Big(\psi^{ms*}_{n^{\prime}}(t)(\sigma_{m^{\prime}n^{\prime}}(\cosh r^{s}_{m^{\prime}}\hat{a}^{in}_{m^{\prime}}+\sinh r^{s}_{m^{\prime}}\hat{a}^{in{\dagger}}_{m^{\prime}})$
$~~~~~~~+\sigma^{*}_{m^{\prime}n^{\prime}}(\cosh r^{s}_{m^{\prime}}\hat{a}^{in{\dagger}}_{m^{\prime}}+\sinh r^{s}_{m^{\prime}}\hat{a}^{in}_{m^{\prime}}))\Big)$
$=\sum_{m,n}e^{2r^{ms}_{n}}(\Re[\sigma_{mn}]e^{r^{s}_{m}}\hat{X}^{in}_{m}-\Im[\sigma_{mn}]e^{-r^{s}_{m}}\hat{P}^{in}_{m})^{2}$
where the integral over $\Delta t$ captures the total time-duration of the pulse, and we have used the orthonormality of modes ($\int dt~\psi_{n}(t)\psi^{*}_{n^{\prime}}(t)=\delta_{nn^{\prime}}$) to insert $\delta_{nn^{\prime}}$ and $\delta_{mm^{\prime}}$ and resolve $m^{\prime}$ and $n^{\prime}$ in the summation.
Bandpass filtering around degeneracy suppresses the imaginary component of $\sigma_{mn}$ as imaginary contributions arise from phase walk-off in the JSI away from degeneracy. Taking $|\Im[\sigma_{mn}]e^{-r_{m}^{s}}|<<|\Re[\sigma_{mn}]e^{r_{m}^{s}}|$, we have:
Figure 9: A) Simulated photon number distributions for unfiltered detection across different pump energies. B) Simulated photon number distributions for 20-nm bandpass filtered detection across different pump energies. C) A plot of the remaining pump energy after pulse propagation (lower remaining energy is a more depleted pump), and the peak histogram bin for the unfiltered and filtered simulations.
$\hat{N}=\sum_{m,n}\frac{e^{2r^{ms}_{n}+2r^{s}_{m}}}{4}\sigma_{mn}(\hat{a}^{in}_{m}+\hat{a}^{in{\dagger}}_{m})^{2},$ (17)
If we compare to the case of no squeezing from the squeezer OPA ($\sigma_{mn}=\delta_{mn},r^{s}_{m}=0~\forall~m$):
$\hat{N}=\sum_{n}\frac{e^{2r^{ms}_{n}}}{4}(\hat{a}^{in}_{n}+\hat{a}^{in\dagger}_{n})^{2}$ (18)
and assume $M$ squeezer modes and $N$ measurement modes, we show that there are now $MN$ modes with effective squeezing parameters $r^{eff}_{m,n}=r^{ms}_{n}+r^{s}_{m}$, whose contributions are weighted by $\sigma_{mn}$. From here, we can return the quadrature representation $\hat{X}=\frac{1}{\sqrt{2}}(\hat{a}+\hat{a}^{\dagger})$ to simplify Eq.17 as:
$\hat{N}=\sum_{m,n}\frac{e^{2r^{ms}_{n}+2r^{s}_{m}}}{8}\sigma_{mn}(\hat{X}_{m}^{in})^{2}$ (19)
In our experiment, we make the assumption that only the first two modes provide significant contributions to the measured statistics as our Schmidt number after filtering is 1.35, and so we truncate $m$ and $n$ to the range [1,2]. Furthermore, we know that $\sigma_{mn}$ where $m$ and $n$ are of different parity are zero. This simplifies Eq.19 to:
$\hat{N}=\frac{e^{2r^{ms}_{1}+2r^{s}_{1}}}{8}\sigma_{11}(\hat{X}_{1}^{in})^{2}+\frac{e^{2r^{ms}_{2}+2r^{s}_{2}}}{8}\sigma_{22}(\hat{X}_{2}^{in})^{2}$ (20)
This allows us to model the measured photon number distribution as the sum of two independent photon number distributions. During the experiment, we treat the fast photodetector as a macroscopic photon number resolving detector such that $N\propto I_{d}$ where $I_{d}$ is the current measured on the detector integrated over a single pulse. For a single mode, the photon number distribution is [30]:
$P^{[1]}_{\langle N\rangle}(N,r)=\frac{1}{\sqrt{2\pi N\langle N\rangle(r)}}e^{-\frac{N}{2\langle N\rangle(r)}}$ (21)
with average photon number $\langle N\rangle(r)=e^{2r}/4$. Since the photon number contributions from each mode are summed at the detector, we can model the total distribution as the convolution of two single-mode distributions:
$P^{[2]}(N,r)=\int_{0}^{\infty}P^{[1]}_{\langle N_{1}\rangle}(N-n,r)P^{[1]}_{\langle N_{2}\rangle}(n,r)dn$ (22)
By fitting Eq.22 to the measured photon number distributions, we isolate the squeezing parameter of the first mode from higher-order modes.
IX Appendix 2: Calculation of the Joint Spectral Intensity and its Modes
Figure 8A shows the JSI of the squeezed vacuum we measure. We calculate our JSI using the dispersive parameters obtained from mode simulations, the input pump pulse energy, and the resources from Ref.[33]. The modes of the JSI are found by performing the Bloch-Messiah decomposition which calculates the modes such that the correlations between pairs of modes are minimized. The first four modes are plotted in Fig. 8B and the fraction of the first eight modes comprising the JSI is shown in Fig. 8C. The JSI decomposition reveals a Schmidt mode number of 7.1. This same calculation can be repeated for the measurement OPA resulting in a Schmidt number of 3.9. Filtering before detection reduces this number to 1.35.
Figure 10: A demultiplexing circuit used to transition from a clock rate $f$ to a clock rate $f/N$.
X Appendix 3: Pump Depletion in Simulation
As discussed in the main text, we model the effects of pump depletion as a shift in the peak of the photon number distribution. Figure 9 shows semi-classical simulation results of vacuum amplification in our OPA generated via a split-step Fourier method. As pump energy increases, the pump pulse depletes more efficiently given the peak-power-enhanced nonlinear interaction, and causes the peak of $P(N)$ to shift away from 0. Fig. 9B shows simulated $P(N)$ with 20-nm bandpass filtering at degeneracy before measurement. In this case, the dominant contribution to the peak shift is pump depletion as multimode effects are suppressed. The difference can be seen in Fig. 9A where in the unfiltered case, peak shifting is more prominent thanks to strong multimode contributions. Peak-shifting as a result of pump depletion has recently been studied in single-mode quantum simulations [37] and experimentally observed [34].
XI Appendix 4: Overcoming the Electronic Bandwidth Limitation
While the OPAs demonstrated in this paper can achieve clock speeds of 6.5 THz, our current experimental setup is still constrained to the speed of the electronics used for photodetection. This problem of needing to return to the electronic domain will continue to exist until we use an all-optical computer to process experimental data [7]. To overcome this, we can employ OPAs as fast switches to demultiplex a clock rate of $f$ down to $f/N$ using $N$ OPAs and photodetectors. Figure 10 shows an example demultiplexer circuit design. An unknown quantum state is amplified and measured inside of the measurement OPA. After measurement, we are left with classical information robust to losses which we split into $N$ channels. Once $N$ clock cycles elapse, we have $N$ pulses in each of the $N$ channels. For each channel, a second pump pulse dubbed the ”gating pump” is sent in to pump the gating amplifier. To achieve demultiplexing, each gating pump is time delayed such that in channel i, the ith pulse will be amplified in the gating amplifier while all other pulses will remain unamplified. While each of the $N$ photodetectors will see a THz stream of pulses, only one out of every $N$ of these pulses will be amplified enough to produce an electronic signal at the photodetector output. Hence, by correctly setting the delays of the gating pump pulses, we can selectively send each pulse to a different photodetectors and effectively lower the system clock rate by a factor of $N$. The feasibility dispersion-engineered OPAs as switches for ultrafast demultiplexing was first studied in [40], however similar schemes have also been presented elsewhere in literature [7].